# Bilevel ZOFO: Bridging Parameter-Efficient and Zeroth-Order Techniques for Efficient LLM Fine-Tuning and Meta-Training

### v1

@@ 4.2 Multi-Task Fine-Tuning Experiments
Method class non_class qa non_qa
$\rightarrow$ class $\rightarrow$ class $\rightarrow$ qa $\rightarrow$ qa
Zero-shot 34.2 34.2 40.2 40.2
Few-shot 34.9 (1.4) 34.9 (1.4) 40.5 (0.3) 40.5 (0.4)
MetaICL 46.4 (1.1) 37.7 (1.7) 45.5 (0.3) 40.2 (0.6)
Ours (Zero-shot) 34.5 34.3 41.8 40.4
Ours(Tuned) 47.1 42.4 43.5 (1.3) 41.9
@@
4.3 Ablative Studies
We perform an ablation study by varying the regularization parameter $\lambda$ (as defined in Equation (5)) and the number of lower-level training steps between each pair of upper-level updates. Figure 2 shows the results. From Figure 2(a), the effect $\lambda$ appears to be non-linear, indicating the need to find an optimal balance. Nontheless, a moderate value like $10$ or $100$ seems to work reasonably well on all tasks. As anticipated, Figure 2(b) demonstrates that performance generally degrades when the total number of upper-level updates is reduced, suggesting there is a trade-off between latency and performance. While more upper-level updates improve results, they also extend the overall training time.
(a) $\lambda$
(b) Number of lower-level training steps
Figure 2: Ablation over $\lambda$ in (5) and the number of lower-level training steps before each upper-level update.
5 Analysis
@@
Lemma 5.4.
If assumption 5.1 holds and $\lambda=1/\epsilon$, assume that $\nabla^{2}F({\mathbf{{\bm{\theta}}}},\cdot)$ is Lipschitz continuous and $({\mathbf{{\bm{\theta}}}},{\mathbf{p}},{\mathbf{s}})$ is an $\epsilon$-stationary point of (4), then $({\mathbf{{\bm{\theta}}}},{\mathbf{s}})$ is an $\epsilon$-stationary point of (2).
The following is the low effective rank assumption from (Malladi et al., 2023). This assumption avoids dimension $d$ in the total complexity. Following (Malladi et al., 2023), we assume here that ${\mathbf{z}}^{k}$ in (7) is sampled from shpere in $\mathbb{R}^{d}$ with radius $\sqrt{d}$ for ease of illustration.
Assumption 5.5.
@@
Theorem 5.6.
If Assumptions 5.1 and 5.5 hold, by setting
$\begin{split}&\eta=\frac{1}{2\ell},\zeta=\frac{1}{2\ell r},\lambda=\frac{1}{\epsilon},B=O(\sigma^{2}\epsilon^{-2}),\\ &\alpha=O(\epsilon\kappa^{-1}(d+d^{\prime})^{-1.5}),T=O\left(\kappa\log(\kappa\epsilon^{-1})\right),\\ &K=O(\kappa r\epsilon^{-2})\end{split}$
@@
This result matches the complexity in previous ZO minimax algorithm in (Wang et al., 2023) but solves our bilevel optimization problem (2) and does not depend on the dimensionality $d$ thanks to the efficient rank assumption 5.5, providing efficiency guarantee for our algorithm.
6 Conclusions
In this work, we introduced a novel bilevel optimization framework designed to mitigate the downsides of PEFT and zeroth-order full model fine-tuning. We propose a new method that is more efficient than existing bilevel methods and thus more suitable for tuning full pre-trained large language models. Theoretically, we provide convergence guarantees for this new method. Empirically, we show that this method outperforms both zeroth-order and FO PEFT methods in single task settings. Additionally, we show this method is effective and efficient when adapted to do multi-task learning. With competitive and even better performance compared to existing meta-training methods, our method offers a significantly cheaper training process.
Impact Statement
This paper introduces an optimization framework that enhances the efficiency of fine-tuning large language models. By reducing computational costs and memory requirements, the approach promotes environmental sustainability and broadens access to advanced AI tools, promoting accessibility in AI development.
Appendix A Method
A.1 Proofs
@@
Lemma A.1.
(Lemma 1.2.3, Theorem 2.1.8 and Theorem 2.1.10 in (Nesterov, 2013))
@@
From lemma A.1 and the definition of $\epsilon$-stationary point (in definition 5.3) we can get the following lemma.
Lemma A.2.
Suppose assumption 5.1 holds and $({\mathbf{x}}_{\epsilon},{\mathbf{y}}_{\epsilon})$ is an $\epsilon$-stationary point of $\min_{\mathbf{x}}\max_{\mathbf{y}}f({\mathbf{x}},{\mathbf{y}})$, let $({\bm{\theta}}_{\epsilon},{\mathbf{p}}_{\epsilon})={\mathbf{x}}_{\epsilon}$ we have
@@
here the first inequality is from Lemma A.1 applied to $-F$ and the second inequality from definition 5.3. ∎
The following is a rephrase of theorem 2 in (Lu & Mei, 2024).
@@
(proof of lemma 5.4) By Lemma A.2 and the value of $\lambda$ we have
@@
Therefore, by Theorem 2 in (Lu & Mei, 2024) we have $\mathbb{E}[\|\nabla F({\bm{\theta}},{\mathbf{p}}^{*}({\bm{\theta}}))\|]\leq O(\epsilon)$ and Lemma 5.4 is proven. ∎
A.1.2 proof of theorem 5.6
Based on Lemma 5.4, it suffices to prove that the algorithm 2 outputs an $\epsilon$-stationary point of $\min_{\mathbf{x}}\max_{\mathbf{y}}f({\mathbf{x}},{\mathbf{y}})$. In this section we will prove this conclusion.
@@
Lemma A.3.
(Lemma C.2 in (Zhang et al., 2024a)) Let ${\mathbf{u}}$ be uniformly sampled from the Euclidean sphere $\sqrt{d}\mathbb{{\mathbf{s}}}^{d-1}$ and ${\mathbf{v}}$ be uniformly sampled from the Euclidean ball $\sqrt{d}\mathbb{B}^{d}=\{{\mathbf{x}}\in\mathbb{R}^{d}\mid\|{\mathbf{x}}\|\leq\sqrt{d}\}$. For any function $f({\mathbf{x}}):\mathbb{R}^{d}\to\mathbb{R}$ and $\alpha>0$, we define its zeroth-order gradient estimator as:
^$\hat{\nabla}f_{\alpha}({\mathbf{x}})=\frac{f({\mathbf{x}}+\alpha{\mathbf{u}})-f({\mathbf{x}}-\alpha{\mathbf{u}})}{2\alpha}{\mathbf{u}},$
@@
Lemma A.4.
If assumption 5.1 holds, for $f_{\alpha}$ defined in Lemma A.3, $\nabla_{\mathbf{x}}f_{\alpha}({\mathbf{x}},{\mathbf{y}})$ is $\ell$-continuous on ${\mathbf{y}}$, i.e.
@@
The first inequality is from Assumption 5.1, second and last inequalities from Lemma A.1 and the equation is from the value of $\eta$.
@@
The following lemma is from Theorem 1 in (Malladi et al., 2023).
Lemma A.6.
If Assumption 5.5 holds, there exists a constant $\gamma={\bm{\theta}}(r)$ such that
@@
Finally, we give the proof for Theorem 5.6. In this part we assume both ${\bm{\theta}}$ and ${\mathbf{p}}$ updates with zeroth order gradient for the convenience of analysis and this does not change the order of the total complexity.
@@
(proof of Theorem 5.6)
From Assumption 5.5, taking expectation conditioning on ${\mathbf{x}}^{k}$ and ${\mathbf{y}}^{k+1}$ we have
@@
Here the last inequality is from Lemma A.3 and Lemma A.4.
@@
where the last inequality is from Assumption 5.1.
On the other hand, from Lemma A.3, by letting $\zeta=\frac{1}{2\ell\gamma}$ we have
@@
Thus, by setting parameters as in Theorem 5.6 we have $\min_{k}\mathbb{E}[\|\nabla_{x}f({\mathbf{x}}^{k},{\mathbf{y}}^{k+1})\|^{2}]\leq\epsilon^{2}.$
@@
similar to Lemma A.5 we have $\mathbb{E}[\|\nabla_{x}f({\mathbf{x}}^{k},{\mathbf{y}}^{k+1})\|^{2}]\leq\epsilon^{2}$ by setting $T=O(\kappa\log(\frac{\kappa}{\epsilon}))$ and $B_{1}=O(\epsilon^{-2}).$
Appendix B Experimental Setup
To recall the proposed Algorithm 2, we present a pipeline of the proposed Algorithm 2 in figure 1.
B.1 Single-Task experiments
Following MeZO (Malladi et al., 2023), we evaluate our approach on a range of classification and multiple-choice tasks. In this setting, training and testing are conducted on the same task.
B.1.1 Tasks
We use the following tasks for evaluating the fine-tuning capabilities of Bilevel-ZOFO in a single-task setting.
BoolQ (Clark et al., 2019):
A yes/no question-answering task where each question is paired with a paragraph that contains the answer.
CB (Wang et al., 2019):
The CommitmentBank task involves determining whether a given sentence in context entails, contradicts, or is neutral to a premise.
COPA (Roemmele et al., 2011):
The Choice of Plausible Alternatives (COPA) task requires selecting the most plausible cause or effect from two alternatives for a given premise.
ReCoRD: (Zhang et al., 2018)
The Reading Comprehension with Commonsense Reasoning Dataset (ReCoRD) is a cloze-style task where models must predict masked-out entities in text based on the surrounding context.
RTE (Wang, 2018):
The Recognizing Textual Entailment (RTE) task involves determining whether a given hypothesis is entailed by a provided premise.
SST2 (Wang, 2018):
The Stanford Sentiment Treebank (SST-2) task focuses on binary sentiment classification of sentences as positive or negative.
WiC (Pilehvar & Camacho-Collados, 2018):
The Word-in-Context (WiC) task involves determining whether the same word is used in the same sense in two different sentences.
WinoGrande (Sakaguchi et al., 2021):
A commonsense reasoning task where the goal is to resolve pronoun references in ambiguous sentences by identifying the correct antecedent.
WSC (Levesque et al., 2012):
The Winograd Schema Challenge (WSC) tests a model’s ability to resolve pronoun references in sentences, requiring commonsense reasoning.
SQuAD (Rajpurkar, 2016):
The Stanford Question Answering Dataset (SQuAD) is a reading comprehension task where models must answer questions based on a given passage of text.
B.1.2 PEFT Variants
We utilize three PEFT techniques—prompt-tuning (Lester et al., 2021), prefix-tuning (Li & Liang, 2021), and LoRA (Hu et al., 2022)—for lower-level training to evaluate bilevel-ZOFO across various conditions and resource constraints.
  1. 1.
    LoRA: For all single-task LoRA experiments, we set $r=8$ and $\alpha=16$.
  2. 2.
    Prefix Tuning: We use 5 prefix tokens across all experiments.
  3. 3.
    Prompt Tuning: We configure 10 soft prompt tokens for every experiment.
B.1.3 Hyperparameter Search
Given resource limitations, we focus on sweeping only the learning rate as the key hyperparameter. For MeZO and first-order PEFT experiments, we explore learning rates from the set $\{1e-2,1e-3,1e-4,1e-5,1e-6\}$. For Bilevel-ZOFO, we sweep both the upper-level and lower-level learning rates: $\text{lr}_{\text{upper}}\in\{1e-4,1e-5,1e-6\}$ and $\text{lr}_{\text{lower}}\in\{1e-2,1e-3,1e-4,1e-5\}$. We perform all experiments in tables 5 and 6 using three random seeds and report the average and standard deviation. We also set $\epsilon=1e-3$, following MeZO (Malladi et al., 2023).
B.1.4 Training

### v2

@@ 4.1.2 Results
4.1.3 Memory Profiling and Wall Clock Time Analysis
Figure 3: Memory consumption of baselines and Bilevel-ZOFO for OPT1.3B (batch size 8, A6000ada 48GB). Bilevel-ZOFO demonstrates memory usage comparable to both baselines.
Task MeZO FO Bilevel-ZOFO
Copa 0.299 0.127 0.135
MultiRC 0.622 0.474 0.502
WSC 0.278 0.120 0.164
Table 4: Wallclock time per step of baselines and Bilevel-ZOFO when fine-tuning OPT1.3B. Values are averaged over 3 runs using a batch size of 8 on a single A6000ada 48GB GPU.
Figure 3 demonstrates the memory profiling of Bilevel-ZOFO, MeZO and First-order prefix tuning on four different tasks. Memory consumption of MeZO and first-order PEFT methods varies across tasks, with one occasionally surpassing the other. Each lower-level update in our method matches that of the corresponding PEFT method. Similarly, each upper-level update requires the greater memory usage between MeZO and PEFT under comparable settings. As a result, the total memory requirement of our method corresponds to the maximum memory usage of the PEFT and MeZO experiments. Nonetheless, as demonstrated in Table 2 and 3 and Figure 1, our method outperforms both PEFT and MeZO on most cases and on average.
We also present a wall-clock time analysis of bilevel-ZOFO compared to the baseline. As shown in Table 4, similar to MeZO [33], we observe that zeroth-order steps exhibit higher latency compared to first-order steps. The results indicate that our bilevel-ZOFO achieves comparable delays to the FO-PEFT method while significantly reducing step duration compared to MeZO. Moreover, as highlighted in Table 2, bilevel-ZOFO outperforms both methods on average.
4.2 Ablations
4.2.1 Effect of Hyper-parameters
We perform an ablation study by varying the regularization parameter $\lambda$ (as defined in Equation (5)) and the number of lower-level training steps between each pair of upper-level updates. Figure 4 shows the results. From Figure 4(a), the effect $\lambda$ appears to be non-linear, indicating the need to find an optimal balance. Nontheless, a moderate value like $10$ or $100$ seems to work reasonably well on all tasks. As anticipated, Figure 4(b) demonstrates that performance generally degrades when the total number of upper-level updates is reduced, suggesting there is a trade-off between latency and performance. While more upper-level updates improve results, they also extend the overall training time. We also analyze different data splits for lower and upper level training. The 1:2 split generally performs well, though effectiveness varies by task. Using a separate upper-level dataset, rather than sharing data across both levels, allows our method to adapt more quickly to new tasks in meta-learning.
(a) $\lambda$
(b) # lower-level training steps
(c) Lower-Upper Data Ratio
Figure 4: Ablation over $\lambda$ in (5), the number of lower-level training steps before each upper-level update, and the ratio of lower/upper data.
4.2.2 Effect of Design Choice
Full-model tuning is only practical via zeroth-order (ZO) optimization due to the high cost of first-order (FO) methods in large models, a core assumption of this work. To address MeZO’s sensitivity and slow convergence while leveraging the strengths of FO PEFT, we propose a hybrid bilevel approach that applies FO to PEFT parameters and ZO to the base model. This section evaluates the benefits of using exact gradients for PEFT. Table 5 compares Bilevel-ZOFO with Bilevel-ZOZO, MeZO, and FO PEFT across four benchmarks, demonstrating the gains from both FO usage and the bilevel structure.
Method BoolQ CB COPA SST2
MeZO 0.6927 0.7767 0.7000 0.8214
FO PEFT 0.7150 0.7142 0.7466 0.8016
Bilevel-ZOZO 0.6280 0.6092 0.7146 0.7633
Two-Stage Pipeline 0.7060 0.6786 0.7433 0.8016
Jointly Optimized 0.7209 0.7500 0.7466 0.8148
Bilevel-ZOFO 0.7367 0.7679 0.7633 0.8335
Table 5: Ablation studies of the effect of different design choices of bilevel-ZOFO as well as gains from the bilevel structure itself.
To ensure that performance gains are not simply due to tuning more parameters, we compare against a Two-Stage Pipeline baseline that is identical to Bilevel-ZOFO in parameter count, memory, and runtime. It applies FO prompt tuning for the same number of steps as Bilevel-ZOFO’s lower level, followed by ZO tuning for the same number of upper-level steps. As shown in Table 5, only the bilevel formulation yields significant improvements, reinforcing the importance of the optimization structure. See Appendix D.2.1 for a detailed discussion.
We also compare with a baseline which jointly optimizes the base model parameters and PEFT parameters together with the same objective function Eq. 1 (Jointly Optimized row in Table  5). While this baseline could slightly outperform MeZO and FO PEFT, it still falls short of our bilevel method, reinforcing our approach’s benefits, although it is more resource consuming and has strictly longer steps due to ZO gradient estimation at every iteration.
4.3 Adopting Bilevel ZOFO to Meta-learning
Following Min et al. [34], we evaluate the performance of bilevel-ZOFO as a fast and efficient meta-learning algorithm. We perform experiments using four of the distinct meta-learning settings: classification-to-classification, non-classification-to-classification, QA-to-QA, and non-QA-to-QA. For instance, in non-classification-to-classification setting, we train on a number of non-classification subtasks and test on a number of distinct classification subtasks. Each of these meta-learning tasks includes a set of training sub-tasks and a different set of test sub-tasks. The sub-tasks are sourced from CROSSFIT [55] and UNIFIEDQA [19], comprising a total of 142 unique sub-tasks. These sub-tasks cover a variety of problems, including text classification and question answering all in English. We use GPT2-Large [40] as the base model for these experiments.
We compare our method against several baseline approaches:
  • MetaICL [34]: A method for meta-learning with in-context learning. MetaICL tunes all the parameters of the base model using the first-order method. In both training and testing, the model is given $k$ demonstration examples, $(a_{1},b_{1}),\dots,(a_{k},b_{k})$}, where $b_{i}$ represents either classification labels or possible answers in question-answering tasks, along with one test example $(a,b)$. The input is formed by concatenating the demonstration examples $a_{1},b_{1},\dots,a_{k},b_{k},a$. The model then computes the conditional probability of each label, and the label with the highest probability is selected as the prediction.
We sample 768 examples from each training sub-task. We train MetaICL in their original setting for 30,000 steps. To train our method, we split the training dataset of each sub-task to two subsets, 256 samples as the development dataset for upper-level updates and 512 samples for lower-level training. For each outer iteration of our method, we randomly sample a subset of 5 training tasks. We perform 10 lower-level updates between each pair of upper-level updates. To keep bilevel-ZOFO as lightweight as possible, unlike MetaICL, we DO NOT include demonstration examples in the inputs. Since bilevel-ZOFO uses significantly less memory and has much faster updates compared to MetaICL, theoretically we are able to train it for many more iterations within the same total training duration as MetaICL. However, due to resource constraints, we only train bilevel-ZOFO for 50,000 iterations. Similar to  [33], we did not observe a plateau in performance for bilevel-ZOFO, indicating that further training can yield additional improvements.
For both ICL and MetaICL, during the testing phase the model is given $k=4$ demonstration examples for each test data point. We don’t use demonstration examples in test samples for bilevel-ZOFO evaluation. We evaluate the zero-shot capabilities of our method as well as the performance of the final model LoRA-tuned for 10 additional iterations on 4 demonstration samples from each class of each test sub-task. Similar to [34], we report Macro-averaged F1 as the evaluation metric. See Appendix D.3 for all training details.
Method Training Inference class non_class qa non_qa
FLOPS Peak Mem(bs=1) Step Duration Tokens / Sample $\rightarrow$ class $\rightarrow$ class $\rightarrow$ qa $\rightarrow$ qa
Zero-shot 0 - - X 34.2 34.2 40.2 40.2
Few-shot 0 - - $\sim 5$X 34.9 (1.4) 34.9 (1.4) 40.5 (0.3) 40.5 (0.4)
MetaICL $1.1354\times 10^{18}$ 32GB 0.48s $\sim 5$X 46.4 (1.1) 37.7 (1.7) 45.5 (0.3) 40.2 (0.6)
Ours (Zero-shot) $1.2485\times 10^{18}$ 12GB 0.18s X 34.5 34.3 41.8 40.4
Ours(Tuned) $1.2493\times 10^{18}$ 12GB 0.19s X 47.1 42.4 43.5 (1.3) 41.9
@@ 4.3 Adopting Bilevel ZOFO to Meta-learning
5 Conclusions
In this work, we introduced a novel bilevel optimization framework designed to mitigate the downsides of PEFT and zeroth-order full model fine-tuning. We propose a new method that is more efficient than existing bilevel methods and thus more suitable for tuning full pre-trained large language models. Bilevel-ZOFO preserves PEFT and ZO-like peak memory, reaches target accuracy in fewer iterations than ZO (yielding 2–4× faster time-to-target despite multi-forward ZO steps), and matches or surpasses FO-PEFT at similar per-step cost—offering a practical accuracy–efficiency trade-off for resource-constrained fine-tuning. Theoretically, we provide convergence guarantees for this new method. Empirically, we show that this method outperforms both zeroth-order and FO PEFT methods in single task settings. Additionally, we show this method is effective and efficient when adapted to do multi-task learning. With competitive and even better performance compared to existing meta-training methods, our method offers a significantly cheaper training process.
6 Acknowledgments
This work was made possible by NSF IIS 2347592, 2348169, DBI 2405416, CCF 2348306, CNS 2347617.
Appendix A Related Work
A.1 Zeroth order in fine tuning LLMs
MeZO [33] is the first work to use Zeroth-Order (ZO) methods to finetune LLMs for downstream tasks. They demonstrate that their method is compatible with both full-parameter tuning and parameter-efficient tuning techniques, such as LoRA and prefix tuning, while being significantly more computationally efficient. Zhang et al. [61] provide a benchmark for ZO optimization in the context of LLM fine-tuning, comparing different ZO optimizers and applying the method to various models. Gautam et al. [10] introduce variance reduction techniques into ZO methods for fine-tuning, improving both stability and convergence. In addition, ZO methods are applied in federated fine-tuning by Qin et al. [39] and Ling et al. [25]. Deng et al. [5] implement ZO optimization for softmax units in LLMs. Guo et al. [14] and Liu et al. [29] explore fine-tuning a minimal subset of LLM parameters using ZO methods by sparsifying gradient approximation or the perturbations used in gradient estimation. Tang et al. [49] investigate the privacy of ZO optimization methods.
A.2 Fine-tuning LLMs for Multitask and Few-Shot Learning
Multi-task learning (MTL) enables a model to handle multiple tasks simultaneously, fostering knowledge transfer between tasks and improving overall efficiency [34, 54]. Typical meta-tuning approaches employ First-Order methods to train autoregressive LLMs on a multitask dataset for various tasks [63, 34, 13]. Zhong et al. [63] apply meta-training to tasks such as hate speech detection, question categorization, topic classification, and sentiment classification. Guo et al. [13] adopt the method from Min et al. [34] for generating stylistic text. While Min et al. [34] focus on enhancing the in-context learning ability of the meta-trained model for multitask learning, Zhong et al. [63] focus on improving zero-shot performance. This approach is particularly valuable in low-resource settings, where collecting large labeled datasets can be costly, as is often the case with medical data. In such environments, few-shot learning—where a model is fine-tuned on a high-resource dataset to quickly adapt to new tasks with minimal data—becomes essential [55]. To address the challenges of multi-task and few-shot learning in natural language processing, several meta fine-tuning methods have been proposed [17, 62, 55, 1]. However, traditional meta fine-tuning approaches, such as MetaICL [34], still require full-model first-order gradient calculations, which become computationally expensive with large language models (LLMs) containing billions of parameters. During training, Min et al. [34] sample a task from the dataset for each iteration to perform in-context learning. In contrast to Zhong et al. [63] and Min et al. [34], our approach uses a bilevel structure: the full LLM is fine-tuned at the upper level, while parameter-efficient fine-tuning (PEFT) models are tuned at the lower level. At test time, we freeze the meta-tuned base model and fine-tune only the PEFT model using a few-shot setup, which is both more cost-effective and efficient. Crucially, Min et al. [34] fine tune the full model with first order methods, while we employ a ZO method in meta-tuning the base model at the upper level. Our approach allows us to bypass the need for backpropagation in the meta-model, significantly reducing computational costs.
A.3 First-order Methods for Bilevel Optimization
Solving a bilevel optimization problem is challenging because the function value in the upper-level objective depends on the optimizer of the lower-level problem. This makes it difficult to compute the gradient of the upper-level objective, also known as the hypergradient. Classical methods require calculating Hessian-vector multiplications to approximate the hypergradient [8, 9, 7, 23, 41, 12, 2, 30]. However, when fine-tuning large language models, this process becomes extremely expensive due to the high computational and memory demands.
Recently, new frameworks for bilevel optimization have been introduced [31, 44, 27, 21, 26, 18, 32]. These methods bypass the need for second-order information by reformulating the bilevel problem as a constrained optimization problem. The constraint is penalized, allowing the problem to be tackled as a minimax problem using only first-order information. These methods significantly reduce computational costs by eliminating the need for second-order information. Nevertheless, when fine tuning LLMs, back propagation for calculating the gradient of an LLM is still too expensive.
Liu et al. [27] and Lu and Mei [31] explore the convergence of their proposed methods to the original bilevel problem, while other approaches only demonstrate convergence to the penalized problem. In this paper, we adapt the method from Lu and Mei [31] to approximate part of the upper-level parameters using a ZO approximation in order to address the challenge posed by the large number of training parameters in large language models. We also provide convergence guarantees for this adapted zeroth-order-first-order method.
Appendix B Theoretical guarantees of Bilevel ZOFO
@@
Lemma B.4.
If assumption B.1 holds and $\lambda=1/\epsilon$, assume that $\nabla^{2}F({\mathbf{{\bm{\theta}}}},\cdot)$ is Lipschitz continuous and $({\mathbf{{\bm{\theta}}}},{\mathbf{p}},{\mathbf{s}})$ is an $\epsilon$-stationary point of (4), then $({\mathbf{{\bm{\theta}}}},{\mathbf{s}})$ is an $\epsilon$-stationary point of (2).
The following is the low effective rank assumption from [33]. This assumption avoids dimension $d$ in the total complexity. Following [33], we assume here that ${\mathbf{z}}^{k}$ in (7) is sampled from shpere in $\mathbb{R}^{d}$ with radius $\sqrt{d}$ for ease of illustration.
Assumption B.5.
@@
Theorem B.6.
If Assumptions B.1 and B.5 hold, by setting
$\vskip-5.0pt\begin{split}&\eta=\frac{1}{2\ell},\zeta=\frac{1}{2\ell r},\lambda=\frac{1}{\epsilon},B=O(\sigma^{2}\epsilon^{-2}),\\ &\alpha=O(\epsilon\kappa^{-1}(d+d^{\prime})^{-1.5}),T=O\left(\kappa\log(\kappa\epsilon^{-1})\right),\\ &K=O(\kappa r\epsilon^{-2})\end{split}$
@@
This result matches the complexity in previous ZO minimax algorithm in [53] but solves our bilevel optimization problem (2) and does not depend on the dimensionality $d$ thanks to the efficient rank assumption B.5, providing efficiency guarantee for our algorithm.
Appendix C Method
C.1 Proofs
@@
Lemma C.1.
(Lemma 1.2.3, Theorem 2.1.8 and Theorem 2.1.10 in [35])
@@
From lemma C.1 and the definition of $\epsilon$-stationary point (in definition B.3) we can get the following lemma.
Lemma C.2.
Suppose assumption B.1 holds and $({\mathbf{x}}_{\epsilon},{\mathbf{y}}_{\epsilon})$ is an $\epsilon$-stationary point of $\min_{\mathbf{x}}\max_{\mathbf{y}}f({\mathbf{x}},{\mathbf{y}})$, let $({\bm{\theta}}_{\epsilon},{\mathbf{p}}_{\epsilon})={\mathbf{x}}_{\epsilon}$ we have
@@
here the first inequality is from Lemma C.1 applied to F-F and the second inequality from definition B.3. ∎
The following is a rephrase of theorem 2 in [31].
@@
(proof of lemma B.4) By Lemma C.2 and the value of $\lambda$ we have
@@
Therefore, by Theorem 2 in [31] we have $\mathbb{E}[\|\nabla F({\bm{\theta}},{\mathbf{p}}^{*}({\bm{\theta}}))\|]\leq O(\epsilon)$ and Lemma B.4 is proven. ∎
C.1.2 proof of theorem B.6
Based on Lemma B.4, it suffices to prove that the algorithm 2 outputs an $\epsilon$-stationary point of $\min_{\mathbf{x}}\max_{\mathbf{y}}f({\mathbf{x}},{\mathbf{y}})$. In this section we will prove this conclusion.
@@
Lemma C.3.
(Lemma C.2 in [58]) Let $\mathbf{u}$} be uniformly sampled from the Euclidean sphere $\sqrt{d}\mathbb{{\mathbf{s}}}^{d-1}$ and $\mathbf{v}$} be uniformly sampled from the Euclidean ball $\sqrt{d}\mathbb{B}^{d}=\{{\mathbf{x}}\in\mathbb{R}^{d}\mid\|{\mathbf{x}}\|\leq\sqrt{d}\}$. For any function $f({\mathbf{x}}):\mathbb{R}^{d}\to\mathbb{R}$ and $\alpha>0$, we define its zeroth-order gradient estimator as:
$\hat{\nabla}f_{\alpha}({\mathbf{x}})=\frac{f({\mathbf{x}}+\alpha{\mathbf{u}})-f({\mathbf{x}}-\alpha{\mathbf{u}})}{2\alpha}{\mathbf{u}},$
@@
Lemma C.4.
If assumption B.1 holds, for $f_{\alpha}$ defined in Lemma C.3, $\nabla_{\mathbf{x}}f_{\alpha}({\mathbf{x}},{\mathbf{y}})$ is $\ell$-continuous on $\mathbf{y}$}, i.e.
@@
The first inequality is from Assumption B.1, second and last inequalities from Lemma C.1 and the equation is from the value of $\eta$.
@@
The following lemma is from Theorem 1 in [33].
Lemma C.6.
If Assumption B.5 holds, there exists a constant $\gamma={\bm{\theta}}(r)$ such that
@@
Finally, we give the proof for Theorem B.6. In this part we assume both $\bm{\theta}$} and $\mathbf{p}$} updates with zeroth order gradient for the convenience of analysis and this does not change the order of the total complexity.
@@
(proof of Theorem B.6)
From Assumption B.5, taking expectation conditioning on ${\mathbf{x}}^{k}$ and ${\mathbf{y}}^{k+1}$ we have
@@
Here the last inequality is from Lemma C.3 and Lemma C.4.
@@
where the last inequality is from Assumption B.1.
On the other hand, from Lemma C.3, by letting $\zeta=\frac{1}{2\ell\gamma}$ we have
@@
Thus, by setting parameters as in Theorem B.6 we have $\min_{k}\mathbb{E}[\|\nabla_{x}f({\mathbf{x}}^{k},{\mathbf{y}}^{k+1})\|^{2}]\leq\epsilon^{2}.$
@@
similar to Lemma C.5 we have $\mathbb{E}[\|\nabla_{x}f({\mathbf{x}}^{k},{\mathbf{y}}^{k+1})\|^{2}]\leq\epsilon^{2}$ by setting $T=O(\kappa\log(\frac{\kappa}{\epsilon}))$ and $B_{1}=O(\epsilon^{-2}).$
Appendix D Experimental Setup
D.1 Single-Task experiments
Following MeZO [33], we evaluate our approach on a range of classification and multiple-choice tasks: BoolQ [3], CB [52], CB [52], COPA [42], ReCoRD: [59],RTE [51], SST2 [51], WiC [37], WinoGrande [43]. In this setting, training and testing are conducted on the same task.
Hyperparameter Search Given resource limitations, we focus on sweeping only the learning rate as the key hyperparameter. For MeZO and first-order PEFT experiments, we explore learning rates from the set $\{1e-2,1e-3,1e-4,1e-5,1e-6\}$. For Bilevel-ZOFO, we sweep both the upper-level and lower-level learning rates: $\text{lr}_{\text{upper}}\in\{1e-4,1e-5,1e-6\}$ and $\text{lr}_{\text{lower}}\in\{1e-2,1e-3,1e-4,1e-5\}$. We perform all experiments in tables 7 and 8 using three random seeds and report the average and standard deviation. We also set $\epsilon=1e-3$, following MeZO [33].
D.1.1 Training